Neptune Company produces toys and other items for use in beach and resort areas. A small, inflatable toy has come onto the market that the company is anxious to produce and sell. The new toy will sell for $3 per unit. Enough capacity exists in the company’s plant to produce 16,000 units of the toy each month. Variable expenses to manufacture and sell one unit would be $1.25, and fixed expenses associated with the toy would total $35,000 per month.

The company's Marketing Department predicts that demand for the new toy will exceed the 16,000 units that the company is able to produce. Additional manufacturing space can be rented from another company at a fixed expense of $1,000 per month. Variable expenses in the rented facility would total $1.40 per unit, due to somewhat less efficient operations than in the main plant.

Required:

1. What is the monthly break-even point for the new toy in unit sales and dollar sales.

2. How many units must be sold each month to attain a target profit of $12,000 per month?

3. If the sales manager receives a bonus of 10 cents for each unit sold in excess of the break-even point, how many units must be sold each month to attain a target profit that equals a 25% return on the monthly investment in fixed expenses?

(For all requirements, do not round intermediate calculations.)
Answer #1

Answer:-

1) Solution: Breakeven point in unit sales: 20081.63 units
Working:
On the first 16000 units
Sales price 3
Variable expenses 1.25
Contribution margin 1.75
Above 16000 units
Sales price 3
Variable expenses 1.4
Contribution margin 1.6
Fixed cost for initial 16000 unts 35,000
Less: CM (16000 units * $1.75) 28,000
Remaining uncovered cost 7,000
Monthly rental for additional space 1,000
Total fixed costs covered by remaining sales 4,082
8000/1.96 = 21,880 units
Breakeven units = 16000+ 4082 = 20081.63
20081.63 * $3 = $60244.9
2) Solution: 27,581.63 units
Working: 12000 / 1.6 = 7500 units
Thus total units = 20081.63 + 7500 = 27,581.63
3) Solution: 110081.6 units
Working: Desired monthly expenses: $35000 + $1000 = 36,000
36000 * 25% = 9000
Unit contribution margin : 1.6 - 0.10 = 1.5
Contribution margin = Target profit / Unit contribution margin = 9000 / 0.10 = 90000 units
20081.63 units + 90000 units = 110081.6 units